# EgoBrain: Synergizing Minds and Eyes For Human Action Understanding

### v1

@@ 6 Conclusion and Discussion
We draw on the metaphor that “the eyes are the windows to the mind” to argue that egocentric video can illuminate neural states that EEG alone cannot. Despite its promise, no dataset or systematic study has yet explored EEG–vision synergy in real-world tasks. We construct EgoBrain, the first action understanding dataset that simultaneously captures first-person video and EEG signals, aiming to advance research on vision-brain signal integration. We further develop Brain-TIM as the first multimodal research baseline on EgoBrain. Experimental results show that combining EEG and visual modalities significantly outperforms single-modality approaches, highlighting the potential of multimodal modeling in complex cognitive scenarios. Notably, we find that EEG signals offer critical complementary information to vision, particularly in scenarios with limited e.g., drinking) or ambiguous (e.g., writing vs. drawing) visual cues.
@@ A.1 Feature Extraction
Before performing action recognition, we first extract features from all modalities using pre-trained encoders. EEG signals are processed through LaBraM-Basejiang:iclr2024 , a model pre-trained on 2,500 hours of masked EEG data that generates 2000-dimensional features per channel. Video frames are encoded using VideoMAEtong:nips2022 , pre-trained on EPIC-KITCHENS-100damen:ijcv21 , which outputs 1024-dimensional features per segment. During feature extraction, we adopt the parameter settings for EEG features as suggested in jiang:iclr2024 , while the parameters for egocentric video features follow those introduced in chalk:cvpr2024 .
@@ A.2 Token Preparation
Visual Features: Within each window in $V$, we uniformly down-sample $K$ frames from their corresponding segment, denoted as $\{v_{1}^{i},\ldots,v_{K}^{i}\}$. The superscript $i$ refers to the $i$-th window, and $t_{i}$ is the starting timestamp of the $i$-th window, corresponding to the time interval $[t_{i},t_{i}+\Delta t)$. The timestamps for each sampled frame are denoted as $\{\tau_{1}^{i},...,\tau_{K}^{i}\}$, where $\tau_{k}^{i}=t_{i}+\frac{2k-1}{2K}\cdot\Delta t$, with $k\in\{1,...,K\}$. This formula ensures that the sampled frames are evenly distributed within the time window and are centered within the window. Each frame is resized to $224\times 224$ and normalized using ImageNetdeng:cvpr2009 statistics. These $K$ frames are then passed through a frozen, pre-trained visual encoder $\phi^{v}$ to produce a window-level feature vector $\mathbf{e}^{v}\in\mathbb{R}^{d^{v}}$, where $d^{v}$ represents the feature dimension. Combining the sliding stride $\delta t$, the full video is encoded into a sequence of window-level feature vectors $\mathcal{E}^{v}=\{\mathbf{e}_{1}^{v},\dots,\mathbf{e}_{N}^{v}\}$, where each $\mathbf{e}_{i}^{v}\in\mathbb{R}^{d^{v}}$.
@@
Token Preparation: After the features from both modalities are obtained, the learnable embedding layers $g^{v}(\cdot)$ and $g^{b}(\cdot)$ are applied to $\mathcal{E}^{v}$ and $\mathcal{E}^{b}$, respectively, to map modality-specific features into a shared $D$-dimensional space. As a result, we obtain the visual feature tokens ~$\tilde{\mathcal{E}}^{v}=\{\tilde{\mathbf{e}}_{i}^{v}\in\mathbb{R}^{D}\}_{i=1}^{N}$ and the EEG feature tokens ~$\tilde{\mathcal{E}}^{b}=\{\tilde{\mathbf{e}}_{i}^{b}\in\mathbb{R}^{D}\}_{i=1}^{N}$ without dimension misalignment for further fusion.
@@ A.3 Transformer Encoder
As shown in the Transformer encoder on the right side of Fig. 3, the input sequence $\mathbf{X}$ is processed by a stack of Transformer encoder layers. Each layer consists of a self-attention mechanism followed by a feedforward network, following the architecture proposed in vaswani:nips2017 . The self-attention mechanism enables the model to capture long-range dependencies across different positions in the sequence, while the feedforward network applies non-linear transformations to the input. Each layer uses residual connections and layer normalization to facilitate gradient flow, with the output passed to the next layer to refine the input sequence $\mathbf{X}$ representations.
@@ A.4 Linear Classification
Appendix B More Visualization
To further illustrate the richness and diversity of the EgoBrain dataset, we provide additional visualizations organized by different subjects (Sec. B.1) and different action categories (Sec. B.2). These examples are intended to help researchers gain deeper insights into the dynamic evolution of multimodal signals across individuals and behaviors within EgoBrain.
B.1 Visualization of Different Subjects
To further illustrate inter-subject variability within the EgoBrain dataset, we present representative initial video frames from all 40 participants in Fig. S1 and Fig. S2. These visualizations clearly demonstrate the abundance of cross-subject samples in EgoBrain, highlighting its strength in capturing individual diversity.
@@ B.1 Visualization of Different Subjects
Figure S1: We provide visualizations of initial video frames from participants: P0001 to P0020.
Figure S2: We provide visualizations of initial video frames from participants: P0021 to P0040.
B.2 Visualization of Different Action Categories
To further demonstrate the richness of action categories in the EgoBrain dataset, we provide visualizations covering all 29 annotated action classes. Fig. S3,  S4,  S5,  S6 and ,  S7 present representative egocentric video frames, highlighting the wide range of daily activities captured in the EgoBrain dataset.
Figure S3: Visualization of selected action categories including “Operate” and “Watch”. The egocentric perspective in each sequence offers intuitive insight into the subject’s ongoing motor behavior.
Figure S4: Visualization of selected action categories including “Play(I)” and “Play(II)”. The egocentric perspective in each sequence offers intuitive insight into the subject’s ongoing motor behavior.
Figure S5: Visualization of selected action categories including “Play(III)” and “Write”. The egocentric perspective in each sequence offers intuitive insight into the subject’s ongoing motor behavior.
Figure S6: Visualization of selected action categories including “Read” and “Draw”. The egocentric perspective in each sequence offers intuitive insight into the subject’s ongoing motor behavior.
Figure S7: Visualization of selected action categories including “Eat” and “Drink”. The egocentric perspective in each sequence offers intuitive insight into the subject’s ongoing motor behavior.

### v2

@@ 6 Conclusion and Discussion
We draw on the metaphor that “the eyes are the windows to the mind” to argue that egocentric video can illuminate neural states that EEG alone cannot. Despite its promise, no dataset or systematic study has yet explored EEG–vision synergy in real‐world tasks. We construct EgoBrain, the first action understanding dataset that simultaneously captures first-person video and EEG signals, aiming to advance research on vision-brain signal integration. We further develop Brain-TIM as the first multimodal research baseline on EgoBrain. Experimental results show that combining EEG and visual modalities significantly outperforms single-modality approaches, highlighting the potential of multimodal modeling in complex cognitive scenarios. Notably, we find that EEG signals offer critical complementary information to vision, particularly in scenarios with limited e.g., drinking) or ambiguous (e.g., writing vs. drawing) visual cues.
@@ A.1 Feature Extraction
Before performing action recognition, we first extract features from all modalities using pre-trained encoders. EEG signals are processed through LaBraMJiang et al. (2024), a model pre-trained on 2,500 hours of masked EEG data that generates 2000-dimensional features per channel. Video frames are encoded using VideoMAETong et al. (2022), pre-trained on EPIC-KITCHENS-100Damen et al. (2022), which outputs 1024-dimensional features per segment. During feature extraction, we adopt the parameter settings for EEG features as suggested in Jiang et al. (2024), while the parameters for egocentric video features follow those introduced in Chalk et al. (2024).
@@ A.2 Token Preparation
Visual Features: Within each window in $V$, we uniformly down-sample $K$ frames from their corresponding segment, denoted as $\{v_{1}^{i},\ldots,v_{K}^{i}\}$. The superscript $i$ refers to the $i$-th window, and $t_{i}$ is the starting timestamp of the $i$-th window, corresponding to the time interval $t_{i},t_{i}+\Delta t)$. The timestamps for each sampled frame are denoted as $\{\tau_{1}^{i},...,\tau_{K}^{i}\}$, where $\tau_{k}^{i}=t_{i}+\frac{2k-1}{2K}\cdot\Delta t$, with $k\in\{1,...,K\}$. This formula ensures that the sampled frames are evenly distributed within the time window and are centered within the window. Each frame is resized to $224\times 224$ and normalized using ImageNetDeng et al. (2009) statistics. These $K$ frames are then passed through a frozen, pre-trained visual encoder $\phi^{v}$ to produce a window-level feature vector $\mathbf{e}^{v}\in\mathbb{R}^{d^{v}}$, where $d^{v}$ represents the feature dimension. Combining the sliding stride $\delta t$, the full video is encoded into a sequence of window-level feature vectors $\mathcal{E}^{v}=\{\mathbf{e}_{1}^{v},\dots,\mathbf{e}_{N}^{v}\}$, where each $\mathbf{e}_{i}^{v}\in\mathbb{R}^{d^{v}}$.
@@
Token Preparation: After the features from both modalities are obtained, the learnable embedding layers $g^{v}(\cdot)$ and $g^{b}(\cdot)$ are applied to $\mathcal{E}^{v}$ and $\mathcal{E}^{b}$, respectively, to map modality-specific features into a shared $D$-dimensional space. As a result, we obtain the visual feature tokens $\tilde{\mathcal{E}}^{v}=\{\tilde{\mathbf{e}}_{i}^{v}\in\mathbb{R}^{D}\}_{i=1}^{N}$ and the EEG feature tokens $\tilde{\mathcal{E}}^{b}=\{\tilde{\mathbf{e}}_{i}^{b}\in\mathbb{R}^{D}\}_{i=1}^{N}$ without dimension misalignment for further fusion.
@@ A.3 Transformer Encoder
As shown in the Transformer encoder on the right side of Fig. 3, the input sequence $\mathbf{X}$ is processed by a stack of Transformer encoder layers. Each layer consists of a self-attention mechanism followed by a feedforward network, following the architecture proposed in Vaswani et al. (2017). The self-attention mechanism enables the model to capture long-range dependencies across different positions in the sequence, while the feedforward network applies non-linear transformations to the input. Each layer uses residual connections and layer normalization to facilitate gradient flow, with the output passed to the next layer to refine the input sequence $\mathbf{X}$ representations.
@@ A.4 Linear Classification
Appendix B EEG Preprocessing Pipeline
Following LaBraM(Jiang et al., 2024)’s preprocessing pipeline to ensure compatibility across modalities, we first applied a band-pass filter of 0.1–75 Hz to suppress low-frequency noise, followed by a 50 Hz notch filter to eliminate power-line interference. The signals were then resampled to 200 Hz and normalized by scaling the EEG amplitudes from their raw range (0.1-0.1 mV to $0.1$ mV) to approximately 1-1 to $1$, with 0.1 mV set as the unit. Unused or noisy channels were removed to further improve signal quality. Finally, raw EEG recordings (e.g., in .edf format) were converted into HDF5 files to facilitate efficient storage and training.
Appendix C More Visualization
To further illustrate the richness and diversity of the EgoBrain dataset, we provide additional visualizations organized by different subjects (Sec. C.1) and different action categories (Sec. C.2). These examples are intended to help researchers gain deeper insights into the dynamic evolution of multimodal signals across individuals and behaviors within EgoBrain.
C.1 Visualization of Different Subjects
To further illustrate inter-subject variability within the EgoBrain dataset, we present representative initial video frames from all 40 participants in Fig. 6 and Fig. 7. These visualizations clearly demonstrate the abundance of cross-subject samples in EgoBrain, highlighting its strength in capturing individual diversity.
@@ C.1 Visualization of Different Subjects
Figure 6: We provide visualizations of initial video frames from participants: P0001 to P0020.
Figure 7: We provide visualizations of initial video frames from participants: P0021 to P0040.
C.2 Visualization of Different Action Categories
To further demonstrate the richness of action categories in the EgoBrain dataset, we provide visualizations covering all 29 annotated action classes. Fig. 8,  9,  10,  11 and ,  12 present representative egocentric video frames, highlighting the wide range of daily activities captured in the EgoBrain dataset.
Figure 8: Visualization of selected action categories including “Operate” and “Watch”. The egocentric perspective in each sequence offers intuitive insight into the subject’s ongoing motor behavior.
Figure 9: Visualization of selected action categories including “Play(I)” and “Play(II)”. The egocentric perspective in each sequence offers intuitive insight into the subject’s ongoing motor behavior.
Figure 10: Visualization of selected action categories including “Play(III)” and “Write”. The egocentric perspective in each sequence offers intuitive insight into the subject’s ongoing motor behavior.
Figure 11: Visualization of selected action categories including “Read” and “Draw”. The egocentric perspective in each sequence offers intuitive insight into the subject’s ongoing motor behavior.
Figure 12: Visualization of selected action categories including “Eat” and “Drink”. The egocentric perspective in each sequence offers intuitive insight into the subject’s ongoing motor behavior.